a bob hanged to astrirg in a box .it is diisplaced to an agle of 45 .then fid the acceleration of the box at this moment.

Grade:12th Pass

5 Answers

Shiva soni
11 years ago

answer is gsin45,that is 10/1.4142.as it would have only tangential accelaration.

mohit yadav
11 years ago

this is the acceleration of the bob.we have to find the acceleration oof the box.

Akash Kumar Dutta
11 years ago

                         /|

                       /x| g

                     /    |

                   /____|

                     a

 

tan x= a/g   

1=a/g

a=g

hence. acc of the block is opposite to the direction of its displacement

|a|=g=10 m/s^2.(ANS).

FITJEE
10 years ago

a=g/4

vishal ojha
10 years ago

gtan(theta)